OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

are you sure both 2 are negative?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

-2x-6=-2x-2-4 -2x-6=-2x-6 0=0 - infinity solutions

OpenStudy (anonymous):

@jhonyy9 negative times a negative is positive

OpenStudy (anonymous):

never mind i thought the 3 was a negative

jhonyy9 (jhonyy9):

Nismo450Z yes this is right --- negativ times a negativ is positive --- yes